# A Tutorial on Beyond-Diagonal Reconfigurable Intelligent Surfaces: Modeling, Architectures, System Design and Optimization, and Applications

### v1

@@ VII-D.2 Future Research Direction
Based on the above discussions, before going to the design of the codebook, it remains unexplored but is important to establish suitable and rigorous criteria to evaluate if a codebook for a specific BD-RIS architecture is “good”. With these criteria, it is interesting to explore the relationship between the weights in a neural network and codewords in a codebook, such that learning-based methods can be adopted to facilitate the codebook design. Beyond this, the idea of end-to-end learning for communication systems [172] can be potentially adopted in practical BD-RIS-aided systems without even knowing the channel model.
VII-E New BD-RIS Architectures
@@
BD-RIS represents a general class of RIS architectures that are more flexible than D-RIS as they include additional tunable components interconnecting the RIS elements with each other. Thanks to this additional flexibility, BD-RIS can further improve the channel shaping capabilities of D-RIS, leading to superior performance in wireless systems, e.g., in terms of channel gain in single-user systems [26] or sum-rate in multi-user systems [36]. Since this performance gain comes at the cost of increased hardware complexity due to the additional impedance components, the fundamental question that arises when considering the deployment of BD-RIS is “Is the additional hardware complexity justified by the resulting performance improvements?
VII-E.1 Challenge
@@
To answer this question, the limits of the trade-off between performance and hardware complexity enabled by BD-RIS need to be studied. In this direction, a model of BD-RIS architectures based on graph theory has been used in previous literature to derive the tree-, forest-, stem-, and band-connected BD-RIS architectures as efficient BD-RIS architectures with reduced complexity [40, 42, 41]. In addition, the set of BD-RIS architectures achieving the Pareto frontier of this performance-complexity trade-off has been characterized for SISO systems, uni-polarized [111] as well as dual-polarized [118]. However, only a limited number of BD-RIS architectures have been explored so far, which are proven to achieve a good trade-off between performance and complexity only under ideal assumptions, e.g., arbitrarily reconfigurable components, lossless RIS, and absence of mutual coupling.
VII-E.2 Future Research Direction
@@
From the graph theoretical model of BD-RIS, we find that the number of possible $M$-element BD-RIS architectures is $2^{M(M-1)/2}$, as there are $M(M-1)/2$ possible interconnections between the $M$ elements. This huge number of potential BD-RIS architectures offers a vast design space that remains largely unexplored. A promising direction for future research is the development of novel BD-RIS architectures that offer favorable performance–complexity trade-offs, particularly under practical constraints such as hardware impairments and real-world non-idealities.
VIII Conclusion
@@
[Preliminaries on Multi-Port Network Analysis] Multi-port network analysis, as a simplified way to understand transmission properties compared to field analysis using Maxwell’s equations, is a powerful and useful technique for modeling and studying wireless systems [34]. Specifically, each antenna in a wireless system is regarded as a port, whose behavior is characterized by its terminal voltage, current, incident wave, and reflected wave. This technique has been used to model and analyze MIMO systems [173, 174, 175], where some hardware impairments, such as mutual coupling and antenna mismatching, can be explicitly captured based on physics-consistent assumptions. Recently, this technique has also been used to model and study RIS-aided wireless communication systems [26, 35, 18, 107], by regarding transmit antennas, receive antennas, and RIS elements as a whole multi-port network. In this appendix, we briefly explain the concept of this technique, together with its three important parameters, namely, impedance, admittance, and scattering parameters, to lay the foundation for the modeling of BD-RIS [34].
Figure 37: An arbitrary $M$-port microwave network.
-A Impedance, Admittance, and Scattering Parameters
Consider an arbitrary $M$-port microwave network, as depicted in Fig. 37. At the $m$-th port, the terminal voltage, current, incident wave, and reflected wave are given by $v_{m}\in\mathbb{C}$, $i_{m}\in\mathbb{C}$, $a_{m}\in\mathbb{C}$, and $b_{m}\in\mathbb{C}$. Define the terminal voltage vector $\mathbf{v}=[v_{1},\ldots,v_{M}]^{\mathsf{T}}\in\mathbb{C}^{M\times 1}$, the current vector $\mathbf{i}=[i_{1},\ldots,i_{M}]^{\mathsf{T}}\in\mathbb{C}^{M\times 1}$, the incident wave vector $\mathbf{a}=[a_{1},\ldots,a_{M}]^{\mathsf{T}}\in\mathbb{C}^{M\times 1}$, and the reflected wave vector $\mathbf{b}=[b_{1},\ldots,b_{M}]^{\mathsf{T}}\in\mathbb{C}^{M\times 1}$ for the $M$-port network. These four vectors are related to each other by
$\mathbf{v}=\mathbf{a}+\mathbf{b},~{}\mathbf{i}=\frac{\mathbf{a}-\mathbf{b}}{Z_{0}}.$ (114)
The impedance matrix $\mathbf{Z}\in\mathbb{C}^{M\times M}$ of the $M$-port network then relates the terminal voltage vector and current vector by $\mathbf{v}=\mathbf{Z}\mathbf{i}$. In the impedance matrix $\mathbf{Z}$, each entry $[\mathbf{Z}]_{m,n}$ is the impedance between ports $m$ and $n$ when all other ports are open-circuited, i.e.,
$[\mathbf{Z}]_{m,n}=\frac{v_{m}}{i_{n}}\Big{|}_{i_{k}=0,\forall k\neq n}.$ (115)
Similarly, the admittance matrix $\mathbf{Y}\in\mathbb{C}^{M\times M}$ of the $M$-port network relates the current vector and voltage vector by $\mathbf{i}=\mathbf{Y}\mathbf{v}$. Therefore, we have $\mathbf{Y}=\mathbf{Z}^{-1}$. In $\mathbf{Y}$, each entry $[\mathbf{Y}]_{m,n}$ is the admittance between ports $m$ and $n$ when all other ports are short-circuited, i.e.,
$[\mathbf{Y}]_{m,n}=\frac{i_{m}}{v_{n}}\Big{|}_{v_{k}=0,\forall k\neq n}.$ (116)
The scattering matrix $\mathbf{S}\in\mathbb{C}^{M\times M}$ relates the incident wave vector and reflected wave vector by $\mathbf{b}=\mathbf{S}\mathbf{a}$. In $\mathbf{S}$, each entry $[\mathbf{S}]_{m,n}$ is the scattering coefficient from port $n$ to port $m$ when all other ports are terminated in matched loads, i.e.,
$[\mathbf{S}]_{m,n}=\frac{b_{m}}{a_{n}}\Big{|}_{a_{k}=0,\forall k\neq n}.$ (117)
Based on the above illustration, the three parameters are related to each other by
$=(Y_{0}\mathbf{I}_{M}+\mathbf{Y})^{-1}(Y_{0}\mathbf{I}_{M}-\mathbf{Y}),$
-B Reciprocal and Lossless Networks
$\mathbf{Z}=\mathbf{Z}^{\mathsf{T}},~{}\mathbf{Y}=\mathbf{Y}^{\mathsf{T}},~{}\mathbf{S}=\mathbf{S}^{\mathsf{T}}.$ (119)
$\Re\{\mathbf{Z}\}=\mathbf{0}_{M\times M},~{}\Re\{\mathbf{Y}\}=\mathbf{0}_{M\times M},~{}\mathbf{\Theta}^{\mathsf{H}}\mathbf{\Theta}=\mathbf{I}_{M}.$ (121)
Note that scattering parameter analysis is convenient for modeling conventional wireless channels [173, 174, 175] and RIS-aided wireless channels [26] since mathematically friendly channel models can be obtained to facilitate further signal processing, and scattering parameters can be measured more easily in practice than the other two. However, impedance and admittance parameter analysis can be more accurate when modeling hardware impairments related to antenna arrays in wireless systems.

### v2

@@ VIII-D.2 Future Research Direction
Based on the above discussions, before going to the design of the codebook, it remains unexplored but is important to establish suitable and rigorous criteria to evaluate if a codebook for a specific BD-RIS architecture is “good”. With these criteria, it is interesting to explore the relationship between the weights in a neural network and codewords in a codebook, such that learning-based methods can be adopted to facilitate the codebook design. Beyond this, the idea of end-to-end learning for communication systems [ait2019model] can be potentially adopted in practical BD-RIS-aided systems without even knowing the channel model.
VIII-E New BD-RIS Architectures
@@
BD-RIS represents a general class of RIS architectures that are more flexible than D-RIS as they include additional tunable components interconnecting the RIS elements with each other. Thanks to this additional flexibility, BD-RIS can further improve the channel shaping capabilities of D-RIS, leading to superior performance in wireless systems, e.g., in terms of channel gain in single-user systems [shen2021] or sum-rate in multi-user systems [li2022beyond]. Since this performance gain comes at the cost of increased hardware complexity due to the additional impedance components, the fundamental question that arises when considering the deployment of BD-RIS is “Is the additional hardware complexity justified by the resulting performance improvements?
VIII-E.1 Challenge
@@
To answer this question, the limits of the trade-off between performance and hardware complexity enabled by BD-RIS need to be studied. In this direction, a model of BD-RIS architectures based on graph theory has been used in previous literature to derive the tree-, forest-, stem-, and band-connected BD-RIS architectures as efficient BD-RIS architectures with reduced complexity [nerini2024beyond, zhou2024novel, wu2025beyond]. In addition, the set of BD-RIS architectures achieving the Pareto frontier of this performance-complexity trade-off has been characterized for SISO systems, uni-polarized [nerini2023pareto] as well as dual-polarized [nerini2025dual]. However, only a limited number of BD-RIS architectures have been explored so far, which are proven to achieve a good trade-off between performance and complexity only under ideal assumptions, e.g., arbitrarily reconfigurable components, lossless RIS, and absence of mutual coupling.
VIII-E.2 Future Research Direction
@@
From the graph theoretical model of BD-RIS, we find that the number of possible $M$-element BD-RIS architectures is $2^{M(M-1)/2}$, as there are M(M1)/2M(M-1)/2 possible interconnections between the $M$ elements. This huge number of potential BD-RIS architectures offers a vast design space that remains largely unexplored. A promising direction for future research is the development of novel BD-RIS architectures that offer favorable performance-complexity trade-offs, particularly under practical constraints such as hardware impairments and real-world non-idealities.
IX Conclusion